# Dongseok Kim    Hyoungsun Choi    Mohamed Jismy Aashik Rasool    Gisung Oh

### v1

@@ Appendix C Implementation and Computational Details
Stage 1 (Subsection 3.1) forms the screened design $X_{S_{0}}$ by retaining the top-$M$ features by global SHAP score. Stage 2 (Subsection 3.2) applies LARS-lasso (lars_path with the lasso option) to the standardized selection-half design on $S_{0}$ and selects the first $K=5$ variables that enter the path. Stage 3 (Subsection 3.3) fits ordinary least squares for $y^{(\mathrm{sur})}$ on the selected features $S$. We compute two sets of regression summaries: (i) a split-sample fit using only the inference half (independent of the selection step), which yields coefficients, standard errors, $t$-statistics, $p$-values, and confidence intervals that are classically valid and can be viewed as a conservative special case of the selective-inference framework in Section 4; and (ii) a naive fit using the full training data, which illustrates the effect of ignoring selection. The global feature-importance table for the $\phi$-test combines the SHAP scores, the selective summaries, and the naive summaries into a single per-feature table and adds a residual row that aggregates the SHAP mass of unselected features.

### v2

@@ Appendix C Implementation and Computational Details
Stage 1 (Subsection 3.1) forms the screened design $X_{S_{0}}$ by retaining the top-$M$ features by global SHAP score. Stage 2 (Subsection 3.2) applies LARS-lasso (lars_path with the lasso option) to the standardized selection-half design on $S_{0}$ and selects the first $K=5$ variables that enter the path. Stage 3 (Subsection 3.3) fits ordinary least squares for $y^{(\mathrm{sur})}$ on the selected features $S$. We compute two sets of regression summaries: (i) a split-sample fit using only the inference half (independent of the selection step), which yields coefficients, standard errors, $t$-statistics, $p$-values, and confidence intervals that are classically valid and can be viewed as a conservative special case of the selective-inference framework in Section 4; and (ii) a naive fit using the full training data, which illustrates the effect of ignoring selection. The global feature-importance table for the $\phi$-test combines the SHAP scores, the selective summaries, and the naive summaries into a single per-feature table into a single per-feature table; for presentation we append a residual row that aggregates the SHAP mass of unselected features.